OpenStudy (anonymous):

2y+x=7 and passing through(-5,-4) in standard form

OpenStudy (aggie12341):

y=-1/2x+7/2

OpenStudy (anonymous):

is that in standard form

OpenStudy (anonymous):

how did u do it and how do i get it in standard form

OpenStudy (aggie12341):

ok well in standard form it would be 1/2x+y=7/2 you just add 1/2 to the left side with y because the way its set up is ax+by=c so x and y needs to be on the same side.